OpenStudy (anonymous):

how to you factor 2x^2+5x-7?

OpenStudy (fibonaccichick666):

that's alot of primes. what ideas do you have?

OpenStudy (wolf1728):

2x(x + 2.5) -7

OpenStudy (anonymous):

i have no idea

OpenStudy (fibonaccichick666):

ok, so, first thing is we want to get rid of that nasty 2

OpenStudy (anonymous):

ok

OpenStudy (fibonaccichick666):

so let's divide it out

OpenStudy (fibonaccichick666):

can you do that?

OpenStudy (popitree):

\[=x(2x+7) - (2x+7)\] =(x-1)(2x+7)

OpenStudy (anonymous):

how did you get that?

OpenStudy (fibonaccichick666):

i was just going to complete the square

OpenStudy (fibonaccichick666):

but popitree's is nicer

OpenStudy (popitree):

the form is ax^2+bx+c; so we need some combination so that we can take common out here a.c = 2*-7; = -14; now as the sign is negative we need two numbers when subtracted will result in b that is 5 here 7-2=5 so basically it is 2x^2+x(7-2)-7 then take the commons